# Scaling Laws in Linear Regression: Compute, Parameters, and Data

### v2

@@ 3 Setup
We use $\mathbf{x}\in\mathbb{H}$ to denote a feature vector, where $\mathbb{H}$ is a finite $d$-dimensional or countably infinite dimensional Hilbert space, and $y\in\mathbb{R}$ to denote its label. In linear regression, we measure the population risk of a parameter $\mathbf{w}\in\mathbb{H}$ by the mean squared error,
@@ 6 Risk bounds under a general spectrum
In this section, we present some general results on the upper and lower bounds of the risk of the output of (SGD). Due to the rotational invariance of the sketched matrix $\mathbf{S}$, without loss of generality, we assume the covariance $\mathbf{H}$ is diagonal with non-increasing diagonal entries. Our main results in Section 4 are directly built on the general bounds introduced here.
@@
almost surely, where $(\lambda_{i})_{i=1}^{d}$ are eigenvalues of $\mathbf{H}$ in non-increasing order, $(\tilde{\lambda}_{i})_{i=1}^{d}$ are eigenvalues of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$ in non-increasing order.
@@ Appendix B Proofs in Section 4
when the stepsize $\gamma\leq c$ for some $a$-dependent constant $c>0.$ Here the hidden constants in the bounds may depend only on $(a,b)$. Combining the bounds on $\mathsf{Approx},\mathsf{Bias},\mathsf{Var}$ and noting
$\mathsf{Var}\eqsim\frac{\min\{M,\ (N_{\mathtt{eff}}\gamma)^{1/a}\}}{N_{\mathtt{eff}}}$ $\lesssim\frac{(N_{\mathtt{eff}}\gamma)^{1/{a}}}{N_{\mathtt{eff}}}\lesssim(N_{\mathtt{eff}}\gamma)^{(1-b)/{a}}\lesssim\mathsf{Bias}+\mathsf{Approx}$
yields Theorem 4.2. Here in the second inequality, we use the assumption $b\leq a$.
@@ C.1 An upper bound
Write the singular value decomposition $\mathbf{H}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$, where $\bm{\Lambda}:=\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots\}$ with $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq 0$ and $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}$. Define $\tilde{\mathbf{S}}:=\mathbf{S}\mathbf{U},\tilde{\mathbf{w}}^{*}:=\mathbf{U}^{\top}{\mathbf{w}^{*}}$. Then by Lemma A.1 the approximation error $\mathsf{Approx}=\mathsf{Approx}(\mathbf{S},\mathbf{H},{\mathbf{w}^{*}})$ satisfies
@@
$=\big{\langle}[\mathbf{H}^{1/2}\mathbf{S}^{\top}\mathbf{A}^{-1}\mathbf{S}\mathbf{H}^{1/2}-\mathbf{I}_{p}]^{\otimes 2},\mathbf{H}^{1/2}{\mathbf{w}^{*}}{\mathbf{w}^{*}}^{\top}\mathbf{H}^{1/2}\big{\rangle}.$
Moreover, for any $k\in[p]$
$\mathbf{H}^{1/2}\mathbf{S}^{\top}\mathbf{A}^{-1}\mathbf{S}\mathbf{H}^{1/2}-\mathbf{I}_{p}$ $=\begin{pmatrix}\mathbf{H}^{1/2}_{{0:k}}\mathbf{S}^{\top}_{{0:k}}\\ \mathbf{H}^{1/2}_{{k:\infty}}\mathbf{S}^{\top}_{{k:\infty}}\end{pmatrix}\mathbf{A}^{-1}\begin{pmatrix}\mathbf{S}_{{k:\infty}}\mathbf{H}^{1/2}_{{k:\infty}}&\mathbf{S}_{{k:\infty}}\mathbf{H}^{1/2}_{{k:\infty}}\end{pmatrix}-\mathbf{I}_{p}$
@@ D.1 An upper bound
Similar to the proof of Lemma C.1, we can without loss of generality assume the covariance matrix $\mathbf{H}=\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots,\lambda_{d}\}$ where $\lambda_{i}\geq\lambda_{j}$ for any $i\geq j$. Let $\mathbf{S}\mathbf{H}^{1/2}=\mathbf{\tilde{U}}\begin{pmatrix}\mathbf{\tilde{\Lambda}}^{1/2}&{\mathbf{0}}\end{pmatrix}\mathbf{\tilde{V}}^{\top}$ be the singular value decomposition of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$, where $\mathbf{\tilde{\Lambda}}:=\operatorname{diag}\{\tilde{\lambda}_{1},\tilde{\lambda}_{2},\ldots,\tilde{\lambda}_{d}\}$ is a diagonal matrix diagonal entries in non-increasing order. Define $\mathbf{A}_{k}:=\mathbf{S}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{S}_{k:\infty}^{\top}$. Then it follows from similar arguments as in Lemma C.1 that $\mathbf{A}_{k}$ is invertible.
@@ D.2 A lower bound
almost surely, where $\mathsf{M_{N}}:=\mathbf{S}\mathbf{H}\mathbf{S}^{\top}(\mathbf{I}-2\gamma\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}}$.
@@
Adopt the notations in the proof of Lemma D.1. By definition of the bias term, we have
@@
$=\langle\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\prod_{t=1}^{N}(\mathbf{I}-\gamma_{t}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}},{\mathbf{v}^{*}}^{\otimes 2}\rangle$
$\geq\langle\mathbf{S}\mathbf{H}\mathbf{S}^{\top}(\mathbf{I}-\sum_{t=1}^{N}\gamma_{t}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}},{\mathbf{v}^{*}}^{\otimes 2}\rangle$
$\geq\langle\mathbf{S}\mathbf{H}\mathbf{S}^{\top}(\mathbf{I}-2\gamma\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}},{\mathbf{v}^{*}}^{\otimes 2}\rangle=:\langle\mathsf{M_{N}},{\mathbf{v}^{*}}^{\otimes 2}\rangle,$ (26)
where the third line uses $\mathbf{I}_{M}-2\gamma_{t}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\succ{\mathbf{0}}_{M}$ for all $t\in[N]$ established in the proof of Lemma D.1, $\sum_{i=1}^{N}\gamma_{i}\leq 2\gamma N_{\mathtt{eff}}$, and the fact that $(1-w)(1-v)\geq 1-w-v$ for $w,v>0$. Substituting the definition of ${\mathbf{v}^{*}}$ in Eq. (5) into the expression, we obtain
@@ F.1 Matching bounds for the average of (SGD) iterates under power-law spectrum
We also verify the observations in Theorem F.6 via simulations. We adopt the same model and setup as in Section 5 but use the average of iterates of fixed stepsize  (SGD) (denoted by $\bar{\mathbf{v}}_{N}$) as the predictor. From Figure 7 and 12 we see that the expected risk ${\mathbb{E}}\mathcal{R}(\bar{\mathbf{v}}_{N})$ also scales following a power-law relation in both sample size $N$ and model size $M$. Moreover, the fitted exponents match our theoretical predictions in Theorem F.6.
Refer to captionRefer to caption
@@ G.1 General concentration results
Let $(\lambda_{i})_{i\geq 1}$ be the eigenvalues of $\mathbf{H}$ in non-increasing order. Let $(\tilde{\lambda}_{i})_{i=1}^{M}$ be the eigenvalues of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$ in non-increasing order. Then there exists a constant $c>1$ such that for every $M\geq 0$ and every $0\leq k\leq M$, with probability $\geq 1-e^{-\Omega(M)}$, we have
@@
With probability at least $1-e^{-\Omega(M)}$, we have
@@
$\lesssim\sqrt{\frac{k+\log(1/\delta))}{M}}.$
So we have, with probability $\geq 1-e^{-\Omega(M)}$,
@@ G.2 Concentration results under power-law spectrum
Let $(\tilde{\lambda}_{i})_{i=1}^{M}$ denote the eigenvalues of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$ in an non-increasing order. Using Lemma G.1 with $k=M/c$ for some sufficiently large constant $c$ and noting that $\sum_{i>k}i^{-{a}}\eqsim k^{1-{a}}$, we have
@@
with probability at least $1-e^{-\Omega(M)}$, where $c_{1},c_{2},c_{3}>0$ are some universal constants. Choosing $k_{0}=M/c^{2}$ for some sufficiently large constant $c>0$, we further obtain
@@ G.3 Concentration results under logarithmic power-law spectrum
Lemma G.6 (Proof of Theorem 6 in (Bartlett et al., 2020)).

### v3

@@ 3 Setup
We denote a feature vector by $\mathbf{x}\in\mathbb{H}$, where $\mathbb{H}$ is a Hilbert space that is either finite-dimensional or countably infinite-dimensional, with dimension $d:=\dim(\mathbb{H})$. The corresponding label is denoted by $y\in\mathbb{R}$. In linear regression, we measure the population risk of a parameter $\mathbf{w}\in\mathbb{H}$ by the mean squared error,
@@ 6 Risk bounds under a general spectrum
In this section, we present some general results on the upper and lower bounds of the risk of the output of (SGD). Due to the rotational invariance of the sketched matrix $\mathbf{S}$, without loss of generality, we assume the covariance $\mathbf{H}\in{\mathbb{R}}^{d\times d}$ is diagonal with non-increasing diagonal entries. Our main results in Section 4 are directly built on the general bounds introduced here.
@@
almost surely, where $(\lambda_{i})_{i=1}^{d}$ are eigenvalues of $\mathbf{H}$ in non-increasing order, $(\tilde{\lambda}_{i})_{i=1}^{M}$ are eigenvalues of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$ in non-increasing order.
@@ Appendix B Proofs in Section 4
when the stepsize $\gamma\leq c$ for some $a$-dependent constant $c>0.$ Here the hidden constants in the bounds may depend only on $(a,b)$. Combining the bounds on $\mathsf{Approx},\mathsf{Bias},\mathsf{Var}$ yields Theorem 4.2.
@@ C.1 An upper bound
Write the singular value decomposition $\mathbf{H}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$, where $\bm{\Lambda}:=\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots,\lambda_{d}\}$ with $\lambda_{1}\geq\lambda_{2}\geq\ldots\lambda_{d}\geq 0$ and $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}$. Define $\tilde{\mathbf{S}}:=\mathbf{S}\mathbf{U},\tilde{\mathbf{w}}^{*}:=\mathbf{U}^{\top}{\mathbf{w}^{*}}$. Then by Lemma A.1 the approximation error $\mathsf{Approx}=\mathsf{Approx}(\mathbf{S},\mathbf{H},{\mathbf{w}^{*}})$ satisfies
@@
$=\big{\langle}[\mathbf{H}^{1/2}\mathbf{S}^{\top}\mathbf{A}^{-1}\mathbf{S}\mathbf{H}^{1/2}-\mathbf{I}_{d}]^{\otimes 2},\mathbf{H}^{1/2}{\mathbf{w}^{*}}{\mathbf{w}^{*}}^{\top}\mathbf{H}^{1/2}\big{\rangle}.$
Moreover, for any $k\in[d]$
$\mathbf{H}^{1/2}\mathbf{S}^{\top}\mathbf{A}^{-1}\mathbf{S}\mathbf{H}^{1/2}-\mathbf{I}_{d}$ $=\begin{pmatrix}\mathbf{H}^{1/2}_{{0:k}}\mathbf{S}^{\top}_{{0:k}}\\ \mathbf{H}^{1/2}_{{k:\infty}}\mathbf{S}^{\top}_{{k:\infty}}\end{pmatrix}\mathbf{A}^{-1}\begin{pmatrix}\mathbf{S}_{{k:\infty}}\mathbf{H}^{1/2}_{{k:\infty}}&\mathbf{S}_{{k:\infty}}\mathbf{H}^{1/2}_{{k:\infty}}\end{pmatrix}-\mathbf{I}_{d}$
@@ D.1 An upper bound
Similar to the proof of Lemma C.1, we can without loss of generality assume the covariance matrix $\mathbf{H}=\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots,\lambda_{d}\}$ where $\lambda_{i}\geq\lambda_{j}$ for any $i\geq j$. Let $\mathbf{S}\mathbf{H}^{1/2}=\mathbf{\tilde{U}}\begin{pmatrix}\mathbf{\tilde{\Lambda}}^{1/2}&{\mathbf{0}}\end{pmatrix}\mathbf{\tilde{V}}^{\top}$ be the singular value decomposition of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$, where $\mathbf{\tilde{\Lambda}}:=\operatorname{diag}\{\tilde{\lambda}_{1},\tilde{\lambda}_{2},\ldots,\tilde{\lambda}_{M}\}$ is a diagonal matrix with diagonal entries in non-increasing order. Define $\mathbf{A}_{k}:=\mathbf{S}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{S}_{k:\infty}^{\top}$. Then it follows from similar arguments as in Lemma C.1 that $\mathbf{A}_{k}$ is invertible.
@@ D.2 A lower bound
almost surely.
@@
Let $\mathsf{M_{N}}:=\mathbf{S}\mathbf{H}\mathbf{S}^{\top}(\mathbf{I}-2\gamma\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}}$. Adopt the notations in the proof of Lemma D.1. By definition of the bias term, we have
@@
$=\langle\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\prod_{t=1}^{N/N_{\mathtt{eff}}}(\mathbf{I}-\gamma_{N_{\mathtt{eff}}t}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}},{\mathbf{v}^{*}}^{\otimes 2}\rangle$
$\geq\langle\mathbf{S}\mathbf{H}\mathbf{S}^{\top}(\mathbf{I}-\sum_{t=1}^{N/N_{\mathtt{eff}}}\gamma_{N_{\mathtt{eff}}t}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}},{\mathbf{v}^{*}}^{\otimes 2}\rangle$
$\geq\langle\mathbf{S}\mathbf{H}\mathbf{S}^{\top}(\mathbf{I}-2\gamma\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{2N_{\mathtt{eff}}},{\mathbf{v}^{*}}^{\otimes 2}\rangle=\langle\mathsf{M_{N}},{\mathbf{v}^{*}}^{\otimes 2}\rangle,$ (26)
where the third line uses $\mathbf{I}_{M}-2\gamma_{t}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\succ{\mathbf{0}}_{M}$ for all $t\in[N]$ established in the proof of Lemma D.1, $\sum_{i=1}^{N/N_{\mathtt{eff}}}\gamma_{N_{\mathtt{eff}}i}\leq 2\gamma$, and the fact that $(1-w)(1-v)\geq 1-w-v$ for $w,v>0$. Substituting the definition of ${\mathbf{v}^{*}}$ in Eq. (5) into the expression, we obtain
@@ F.1 Matching bounds for the average of (SGD) iterates under power-law spectrum
We also verify the observations in Theorem F.6 via simulations. We adopt the same model and setup as in Section 5 but use the average of iterates of fixed stepsize  (SGD) (denoted by $\bar{\mathbf{v}}_{N}$) as the predictor. From Figure 3 and 4 we see that the expected risk ${\mathbb{E}}\mathcal{R}(\bar{\mathbf{v}}_{N})$ also scales following a power-law relation in both sample size $N$ and model size $M$. Moreover, the fitted exponents match our theoretical predictions in Theorem F.6.
@@ G.1 General concentration results
Let $(\lambda_{i})_{i\geq 1}$ be the eigenvalues of $\mathbf{H}$ in non-increasing order. Let $(\tilde{\lambda}_{i})_{i=1}^{M}$ be the eigenvalues of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$ in non-increasing order. Then there exists a constant $c>1$ such that for every $M\geq 0$ and every $0\leq k\leq M$, with probability $\geq 1-e^{-\Omega(M)}-e^{-\Omega(k)}$, we have
@@
With probability at least $1-e^{-\Omega(M)}-e^{-\Omega(k)}$, we have
@@
$\lesssim\sqrt{\frac{k+\log(1/\delta)}{M}}.$
So we have, with probability $\geq 1-e^{-\Omega(k)}$,
@@ G.2 Concentration results under power-law spectrum
Let $(\tilde{\lambda}_{i})_{i=1}^{M}$ denote the eigenvalues of $\mathbf{S}\mathbf{H}\mathbf{S}^{\top}$ in non-increasing order. Using Lemma G.1 with $k=M/c$ for some sufficiently large constant $c$ and noting that $\sum_{i>k}i^{-{a}}\eqsim k^{1-{a}}$, we have
@@
with probability at least $1-e^{-\Omega(M)}-e^{-\Omega(k_{0})}$, where $c_{1},c_{2},c_{3}>0$ are some universal constants. Choosing $k_{0}=M/c^{2}$ for some sufficiently large constant $c>0$, we further obtain
@@ G.3 Concentration results under logarithmic power-law spectrum
Lemma G.6 (Theorem 6 in (Bartlett et al., 2020)).